It has been known that a fixed dose of a certain drug results to an average increase of pulse rate by at least 12 beats per minute with a standard deviation of 5. A group of 20 patients given the same dose showed the following increases: 15, 12, 10, 8, 14, 15, 16, 11, 7, 13, 9, 10, 12, 11, 9, 10, 17, 14, 15, 7. Is there proof to show that this group has a lower average increase of pulse rate than the ones in general? Use 0.05 level of significance.


1
Expert's answer
2021-11-19T11:06:15-0500

"n=20 \\

\bar{x} = \frac{15+12+...+15+7}{20}= 11.75 \\

\sigma = 5.0 \\

H_0: \mu ≥12 \\

H_1: \mu < 12"

Test-statistic

"Z = \frac{\bar{x} -\mu}{ \sigma / \sqrt{n}} \\

Z = \frac{11.75-12.0}{5.0/ \sqrt{20}} = -0.22"

P-value = P(Z< -0.22) = 0.4129

(from normal distribution z -table)

Since P-value is greater than significance level α = 0.05 ,we failed to reject the null hypothesis.

Decision: fail to reject H0.

Conclusion :

There is not sufficient evidence to conclude that group has a lower average increase of pulse rate than the ones in general.